OpenStudy (anonymous):

Use the substitution u = tan(x) to evaluate the following. int_0^(pi\/6) (text(tan) ^2 x text( sec) ^4 x) text( ) dx

OpenStudy (freckles):

\[\int\limits_0^\frac{\pi}{6} \tan^2(x)\sec^4(x) dx\]

OpenStudy (freckles):

hints: Pythagorean identity